. Calculate the [OH-] for a solution with a pH value of 10.28
1
Expert's answer
2020-05-26T14:20:54-0400

pH+pOH=14

pOH = 14 − pH

pOH = 14 − 10.28 = 3.72

pOH=−log[OH]

3.72 =−log[OH]

10−(3.72) = [OH]

[OH] = 10−(3.72) = 0.00019 mol/L = 1.9 × 10−4 mol/L

Answer: 1.9 × 10−4 mol/L